IP Nas for the vector field (=.=) - deye where is the surface of region shown in the below
Answer #1

Aus) F.NDS geven vector field - 4xê +42h+4x4 3k wheres is z=4-2/22 +42 using divergence theosam F. Nds = SS. s divf du VE fivश्रा -2 : 168 +4x2 + 982278 - 283%nefite या (-232+ 100 +3-83579 = Fir to था 32+ 32 +64%re - Birg) to 3 श्रा - 64 3 + 555 भा t